# Comprehensive Review of Deep Unfolding Techniques for Next-Generation Wireless Communication Systems

### v2

@@ X Physical Layer Security
Jamming attacks pose a serious threat to the reliable operation of wireless communication systems [249, 250]. A jamming attacker deliberately transmits interference or noise on the same frequencies as legitimate users. This disrupts communication channels, reduces signal quality, and can block data transmission entirely. The interference makes it harder for users to decode messages, leading to connection losses or forced retransmissions. As a result, network efficiency decreases, and latency increases. Marti et al. in [251] explored solutions to combat smart jamming in massive multi-user MIMO systems. In these systems, jamming attacks can selectively interfere at specific times to avoid detection. The authors proposed an optimization-based approach for joint jammer estimation, channel estimation, and data detection, using the BS’s multi-antenna configuration. This method does not require prior knowledge of the jammer’s behavior, making communication more resilient even during sporadic or bursty jamming. Using a deep unfolding approach, the authors developed two iterative algorithms. The first is mitigation estimation and detection (MAED), which uses forward-backward splitting for estimation. The second is soft-output MAED (SO-MAED), an enhanced version that applies deep unfolding to optimize parameters. SO-MAED improves detection accuracy, especially in high-interference environments, and does not require knowledge of the attack type. Error precision and gradient step sizes are key parameters for improving learning stability.
XI Joint Techniques
This section reviews different joint techniques that are aided by deep unfolding method.
XI-A Joint Detection and Channel Estimation
An iterative method for estimating joint channel parameters and selecting symbols using the EM algorithm is presented in [252]. This method performs symbol detection in the presence of Gaussian noise and intersymbol interference in a dispersive channel. The authors model the missing data using a Markov chain, and channel parameter estimation follows the ML criterion. The Markovian properties of the channel state sequence allow for calculating the required likelihoods using the forward-backward algorithm. In [253], the authors explore a V-BLAST-type combination of OFDM with MIMO to improve SE and multi-user downlink throughput. They propose a joint data detection and channel estimation algorithm for MIMO-OFDM that integrates the QR decomposition M-algorithm and Kalman filter. The M-algorithm, combined with QR decomposition, reduces the computational complexity of tree-based search methods [254].
As mentioned in [6], the data-driven approach uses a fully connected network, where each sub-block of the OFDM transceiver is designed. Gao et al. [255] proposed a deep unfolding approach to address the issue of DNN performance being dependent on large training datasets. In contrast to the approach in [6], where a DL architecture completely replaces the conventional modules, the ComNet architecture shown in Fig. 9 incorporates both channel estimation and detection modules. Both modules use conventional communication solutions as inputs to the network. The DL network then refines these coarse inputs. For example, the LS channel estimate, $\bf{\hat{h}_{\text{LS}}}$, is refined using pilot symbols ${\bf{x}}_{p}$ and corresponding received symbols ${\bf{y}}_{p}$ as inputs to the LS sub-network within the channel estimation module. The channel estimate, $\bf{\hat{h}_{\text{LS}}}$, is passed from the LS sub-network to the LS Refine Net sub-network to generate an accurate channel estimate, $\bf{\hat{h}}$. A single-layer DNN is used to design the LS Refine Net. The refined channel estimate, $\bf{\hat{h}}$, from the LS Refine Net is then passed as input to the ZF detector sub-network in the signal detection module. The ZF detector sub-network is also fed with the received data vector, $\bf{y_{\text{D}}}$, and the LS channel estimate, $\bf{\hat{h}_{\text{LS}}}$. This is followed by the ZF Refine Net. The refined channel estimate, $\bf{\hat{h}}$, the received data vector, $\bf{y_{\text{D}}}$, and the ZF estimate, $\bf{\hat{x}_{\text{ZF}}}$, from the ZF detector sub-network are passed to the ZF Refine Net to estimate the transmit binary data, $\hat{b}$. The short dotted line from the LS sub-network to the ZF detector sub-network in Fig. 9 provides a robust candidate for the recovered binary symbols, $\hat{b}^{{}^{\prime}}$, at deQAM. Both the channel estimation and detection modules are trained using an $\ell_{2}$ loss between the estimated binary data, $\hat{b}$, and the raw transmit binary symbol, $b$. Simulations have shown that this design outperforms both LMMSE and FC-DNN [6].
Figure 9: ComNet architecture [255].
An artificial intelligence-aided receiver for a cyclic prefix-free OFDM system with joint detection and channel estimation is explored in [256]. Zhang et al. [257] unfolded the OAMP algorithm for generalized frequency division multiplexing (GFDM) systems with low-resolution analog-to-digital converters. In [258], a general EM algorithm is derived and unfolded to address the impact of channel estimation errors on signal detection. The expectation step estimates the CSI, while the maximization step detects the transmitted signal using the estimated CSI. A modified, trainable projected gradient detector is used in the maximization step. In [259], an ADMM-based unfolded algorithm is proposed for robust MIMO detection, incorporating channel estimation errors. To reduce error propagation due to channel aging, a robust data-aided Kalman filter-based channel tracking method is developed. The pre-estimated data symbols are used to improve channel estimation accuracy, enhancing robust detection performance.
Recently, the energy harvesting technique has gained significant attention as a promising feature for green communication networks. Ambient backscatter communication is an energy-efficient technology that allows devices to transmit data by reflecting existing ambient signals, such as those from Wi-Fi routers or cellular networks, instead of generating their own radio waves. A backscatter device modulates the signal over the ambient radio frequency without an active radio frequency component. However, the backscatter link is weaker than the direct link, which limits its coverage and data rate [260]. In [261], Wang et al. explored an IRS-based ambient backscatter communication system. The proposed network consists of two parts. The first subnetwork calculates the IRS phase shift matrix for binary symbols. The phase shift matrix weights, or phase angle vectors, are updated using a DNN training tool. These updated phase angle vectors are then used to compute the received signal, which is passed to the second subnetwork, the detection module. The EM algorithm is unfolded for symbol detection.
XI-B Joint Activity Detection and Channel Estimation
In [262], Ke et al. addressed massive access in massive MIMO systems by introducing an adaptive scheme for active user detection and channel estimation using compressive sensing techniques. This scheme takes advantage of the sparse traffic patterns of many user devices and the sparse nature of massive MIMO channels in the virtual angular domain. The access latency is reduced due to the use of non-orthogonal pseudo-random pilots for uplink communication. The problem of detecting active users and estimating their channels is formulated as a generalized multiple measurement vector compressive sensing problem. A generalized multiple measurement vector AMP (GMMV-AMP) algorithm is designed to leverage the structured sparsity in the uplink channel matrix. Additionally, a Turbo-GMMV-AMP algorithm is developed to exploit channel sparsity in both the spatial and angular domains. In [263], an EP-based joint active user detection and channel estimation technique for mMTC networks is proposed. The EP algorithm iteratively minimizes the Kullback-Leibler divergence between the target and approximate distributions via moment matching, efficiently finding a tractable close approximation of the target distribution.
Grant-free random access is used for active device detection at the BS to address the latency issue in massive access. The active device detection is formulated as a sparse signal recovery problem. To improve the performance of massive device detection, Qiang et al. [43] proposed a deep unfolded network called AMP-Net, as shown in Fig. 10. The basic principle of AMP-Net is based on the AMP algorithm. The input variables passed through the $T$ layers include the state evolution parameter, $\tau^{2}$, the device state matrix $\bf{X}$, the received signal $\bf{Y}$ at the BS, and the residual matrix $\bf{R}$. The pilot vectors are denoted by $\bf{a_{\text{n}}}$. Each layer contains an MMSE estimator, $\eta_{t}(\cdot)$. The set of trainable parameters consists of four scalar values. The parameters $\lambda_{n}^{t}$ and $\gamma_{n}^{t}$ estimate the activity probability and channel variance, respectively. The parameter $\theta_{n}^{t}$ regulates the denoiser, while $\mu_{n}^{t}$ controls the state evolution parameter, helping facilitate an accurate prediction of the detected target.
Figure 10: The structure of AMP-Net [43].
Traditional orthogonal multiple access fails to achieve massive connectivity for mMTC systems. Random and sporadic data traffic means that only a small fraction of users access the BS during uplink transmission. The joint channel estimation and active user detection problem for uplink mMTC is investigated in [264]. The authors developed a pilot transmission scheme for nonorthogonal identification. Due to the sporadic activity of users, where only a small fraction are active at any time, traditional orthogonal pilots are inefficient. Instead, nonorthogonal pilots are used, and the joint estimation and detection problem is formulated as a compressive sensing problem. The work considers single and multiple antenna configurations at the base station, modeling them as single measurement vector and multiple measurement vector problems, respectively. The authors designed a deep unfolding model based on the ADMM algorithm. For ADMM, the trainable parameters include: (1) an auxiliary variable to compute matrix inversions, (2) a regularization parameter that controls the penalty term in the ADMM optimization, influencing the balance between sparsity and accuracy in the recovered signal, and (3) a threshold parameter in the shrinkage function that determines the sparsity of the output in each iteration. In [265], two deep unfolded neural network architectures are proposed for sporadic user detection and channel estimation in mMTC. The first architecture unfolds a linearized ADMM, applied to a constrained convex problem. Linearized ADMM addresses computationally intractable block coordinate sub-problems by executing a single iteration of the proximal gradient method for each subproblem instead of solving them explicitly [156]. The second architecture unfolds a vector AMP with a denoiser based on ISTA. This unfolded network reduces complexity and improves accuracy compared to the original algorithm.
In [266], a deep unfolded algorithm is proposed for the joint design of compression modules, quantization modules at the APs, and the decompression and detection modules at the CPU for cell-free massive MIMO systems with capacity-limited fronthauls. The objective is to address activity detection in an end-to-end learning framework. Ma et al. [267] investigated joint device activity and data detection in mMTC. They proposed a one-phase non-coherent scheme, where data bits are embedded in the pilot sequences. The BS simultaneously detects active devices and their embedded data bits without explicit channel estimation. The AMP algorithm is unfolded to enhance detection performance by effectively exploiting the pilot activity correlation. In [268], a contention-based asynchronous random access scheme with data length diversity is considered for mMTC. The paper unfolds AMP with backpropagation to address the composite problem of activity detection, channel estimation, and data recovery. Zou et al. [269] formulated the joint activity detection and channel estimation problem as a group-row-sparse matrix recovery problem. A proximal gradient-based unfolding NN is developed to parameterize the algorithmic iterations and solve this problem. In [270], compressive sensing recovery algorithms are used to address the joint active device detection and channel estimation problem for uplink transmission in grant-free based vehicular networks. The ADMM algorithm is unfolded to improve convergence rate and recovery accuracy. Shi et al. [271] considered grant-free massive access in an IoT network. A large number of single-antenna IoT devices sporadically transmit non-orthogonal sequences to a multi-antenna BS. An unfolded DNN framework is developed using the multidimensional shrinkage thresholding operator inherited from the ISTA structure. This network simultaneously achieves low computational complexity and high robustness for solving the joint activity detection and channel estimation problem.
Cell-free technology for massive MIMO enhances mmWave signal coverage [272]. Using deep unfolding techniques, the authors in [273] developed a grant-free access protocol based on mmWave cell-free massive MIMO. The proposed hybrid structure leverages learned vector AMP to reduce the need for prior channel information and achieve faster convergence.
XI-C Joint Channel Estimation and Precoder Design
A codebook design for channel estimation in mmWave communications with a hybrid precoding structure is studied in [274]. To evaluate the performance of a codeword, the authors propose a metric called generalized detection probability (GDP), which considers both the beam pattern’s flatness and the maximum transmission power. Yuan et al. [275] introduced a two-stage wideband hybrid beamforming scheme for THz wireless systems. The model features a multi-antenna BS communicating with a multi-antenna user over frequency-selective fading. Additionally, a probabilistic robust hybrid beamforming scheme is designed to mitigate channel estimation errors.
(a) Long-term
(b) Short-term
Figure 11: JCEHB architecture [276].
Instead of modeling and optimizing hybrid beamforming and channel estimation separately, Kang et al. [276] proposed the joint channel estimation hybrid beamforming (JCEHB) algorithm. The sum rate is maximized for TDD massive MIMO system in a mixed timescale framework. The algorithm’s architecture is shown in Fig. 11. Observe from Fig. LABEL:sub@fig:fig1, the deep-unfolding NN architecture consists of two parts : channel estimation and hybrid beamforming. For channel estimation, the pilot training module is designed by unfolding the recursive least squares (RLS) algorithm. The SSCA algorithm is unfolded to replace the high complexity computation operations in hybrid beamforming. The long-term analog beamforming matrices, are optimized offline with CSI statistics that is constant over long time slots or a superframe. After fixing the analog beamformers, the deep-unfolding NN for channel estimation is trained at each time slot to optimize the digital beamformers, as shown in Fig. LABEL:sub@fig:fig2. In contrast to long-term analog beamforming, short-term digital beamforming matrices are optimized every time slot using real-time low-dimensional effective CSI. Analog beamforming matrices are optimized offline using CSI statistics, while digital beamforming matrices are updated within each time slot based on estimated low-dimensional CSI. The authors claimed that the hybrid approach in a mixed (long-term and short term) timescale framework reduce signaling overhead as compared to separate design of NN in a single timescale.
An unfolded WMMSE-based NN for massive multi-user MIMO under imperfect CSI is studied in [277]. The proposed network enhances performance with low complexity by leveraging CSI statistics for near-optimal precoder solutions, accelerating convergence. In [278], an unfolded sparse Bayesian learning algorithm is derived to capture channel sparsity in a mmWave massive MIMO system with hybrid beamforming. Similarly, [279] proposes a learning-based channel semantic acquisition and beamforming solution for cell-free massive MIMO. A multi-layer perceptron estimates the channel, while a successive over-relaxation-based linear beamforming scheme reduces estimation errors in hybrid beamforming. Wideband precoding under imperfect CSI is discussed in [280]. A robust frequency-selective precoding algorithm maximizes the expected weighted sum rate while considering per-antenna power constraints. This algorithm is unfolded to reduce computational complexity. To improve beamspace channel estimation, [281] introduces a deep unfolding algorithm using the top $\ell_{K,2}$ norm. This norm, a variation of the traditional $\ell_{2}$ norm, focuses on the largest elements in the sparse beamspace channel vector [282]. It helps determine sparsity levels, reducing noise and computational complexity. The iterative trimmed-ridge regression algorithm is unfolded to allow parallel computation. These algorithms are applicable to both narrowband and wideband massive MIMO systems, ensuring versatility in channel estimation.
XI-D Joint Detection and Decoding
An adaptive Bayesian receiver using sequential Monte Carlo for signal detection and decoding in fading channels with known channel statistics is proposed in [283]. The transmitted signal is treated as missing data, and multiple signal samples are sequentially estimated based on observed signals. These estimated sequences approximate the Bayesian estimate of both transmitted signals and channel states. Adaptive receiver algorithms are developed for both uncoded and convolutionally coded systems. El Gamal et al. [284] present an iterative receiver for joint detection and decoding of CDMA signals. The multiuser detector follows the MMSE criterion. A modified symbol-aided demodulation technique extracts soft information from previous decoding iterations to improve channel estimates, enhancing joint detection and decoding efficiency.
Traditionally, channel estimation, signal detection, and channel decoding in a MIMO receiver are performed sequentially and independently. This disjoint approach leads to suboptimal performance. To address this, Sun et al. [285] propose receivers that integrate these functions, reducing error propagation and processing delay. Their design targets MIMO ultra-reliable low-latency communication (URLLC) systems. The model uses short channel codes and a limited number of pilot symbols for channel estimation. The authors employ the ADMM algorithm to solve large-scale non-convex optimization problems, deriving closed-form solutions for each iteration. Two NNs, JCDDNet-G for Gaussian channels and JCDDNet-S for sparse mmWave channels, are built by unfolding the ADMM algorithms as shown in Fig. 12. Simulations show that the proposed trainable JCDD receivers significantly outperform traditional turbo receivers.
Figure 12: JCDDNet-S/G architecture [285].
In [286], Zhang et al. propose a receiver framework based on a deep unfolded NN structure. The joint signal detection and channel decoding modules are modeled as multi-layer deep feed-forward networks. Signal detection is achieved by unfolding the EP algorithm, while channel decoding is based on unfolded turbo decoding. Building on this work, Zhang et al. develop an unfolding-based framework for MIMO detectors in [287]. This framework adaptively determines trainable parameters, allowing the detector to adjust effectively to varying channel conditions.
@@ XI-D Joint Detection and Decoding
XII Challenges in Developing Deep Unfolding Models and Future Research Scopes
@@
XII-A Limitations of Deep Unfolding Techniques
Although deep unfolding has found applications in various design problems, it has limitations. Most of these limitations arise from the fact that deep unfolding is not an independent design methodology; it must be built upon existing conventional optimization frameworks. Consequently, developing deep unfolding solutions faces the following challenges:
(C1) Contrasts between conventional optimization and deep unfolding: There is a notable contradiction between the development methodologies of conventional optimizers and deep unfolding models. Most conventional solutions for transceiver design and signal processing in wireless communications are based on convex optimization techniques. However, many wireless design problems are inherently non-convex and highly coupled due to factors such as fractional SINR, constant-modulus constraints (e.g., for analog beamforming), and mixed-integer variables (e.g., in symbol detection or antenna selection). These challenges are typically addressed using a two-layer approach. In the first layer, design variables are decoupled through alternating optimization (AO) or block coordinate ascent/descent, breaking the original problem into smaller sub-problems, each focusing on a subset of variables that are solved alternately. In the second layer, techniques such as SCA or the ADMM are applied to solve each sub-problem, often introducing additional variables. Although this two-layer approach transforms the original non-convex problem into a series of convex ones that can be solved using convex optimization tools like CVX, it significantly increases the number of variables and sub-problems and leads to nested loops in the solution process. These loops, corresponding to the two layers, must iterate until convergence, adding considerable complexity. In contrast, deep unfolding models typically approach the problem as a unified whole, solving all variables in parallel, facilitating the development of unfolding architectures. This fundamental difference makes it challenging to develop suitable algorithms for unfolding. Additionally, the increased complexity in variables and sub-problems inherent in conventional methods further complicates the design of deep unfolding frameworks, directly contributing to their overall complexity.
(C2) Challenges in unfolding principle algorithms: While deep unfolding typically enhances conventional iterative algorithms, not all can be effectively unfolded. This limitation arises from the core principle of deep unfolding, which structures the iterations of an iterative algorithm as layers within a DNN. Complex operations like matrix inversion, matrix decomposition, or the use of convex optimization tools such as CVX or YALMIP (with solvers like MOSEK, SeDuMi, etc.) are difficult to implement in a DNN framework. In large-scale systems such as massive MIMO, even matrix multiplications result in high complexity for unfolding models.
@@ XII-A Limitations of Deep Unfolding Techniques
(C3) Challenge in handling design constraints and objectives: One of the biggest challenges in developing deep unfolding models lies in managing the design constraints formulated in the problem. These constraints require the DNN layers’ outputs to meet specific requirements, such as total power, QoS, or other element-wise constraints. Simple constraints, like power or constant modulus, can often be enforced through normalization [118]. For instance, the $tanh$ activation function can be employed to impose constant-modulus constraints on analog beamforming coefficients or sensing waveforms, similar to black-box DL architectures. However, QoS constraints are significantly more complex. While they resemble activation functions, none of the existing activation functions can adequately manage such intricate requirements. One approach is to incorporate these constraints into the design objective, but this method often results in a problem that is not equivalent to the original constrained one, making the solution potentially infeasible. Properly modeling such constraints within a deep unfolding DNN remains highly challenging, hindering the development of more advanced deep unfolding models.
@@
Not only do complex constraints pose challenges for developing deep unfolding models, but intricate objective functions also present difficulties. Generally, deep unfolding models are trained based on the objective function. For instance, unfolding architectures like DetNet [49], FSNet [152], and ScNet [288] for symbol detection are trained to minimize the Euclidean distance between the model’s output (symbol vector) and the true transmitted symbol vector. When labels are unavailable, as in beamforming problems, unfolding networks use the objective function as the loss function for minimization or its negative for maximization problems. Typical deep unfolding techniques [84, 75] require calculating the gradient of the objective function. However, this becomes challenging when the objective function is a complex, high-order, or fractional function of the design variables, especially when those variables are vectors or matrices. Therefore, more dynamic methodologies are needed for developing unfolding models that can handle such intricate functions efficiently.
@@
(C4) Need for black-box models in sophisticated mappings: While deep unfolding is powerful, certain inference tasks still require black-box DNN models, especially when there is no clear relationship between the input data, output solution, and performance metrics. For instance, in multi-modal sensing-assisted beam tracking, tasks may involve complex mappings from environmental data—such as sensory information from sensors and cameras—to the optimal beam [289, 290]. In such cases, there is no explicit relationship between environmental images and the optimal beam, and standard communication metrics like throughput or SINR cannot be directly expressed as functions of these images. These problems cannot be formulated as optimization tasks solvable with principle algorithms, making it difficult to develop unfolding models that map environmental information to beam patterns or blockage states.
@@
The above three challenges C1C4 also outline the scenarios wherein deep unfolding models cannot be built or does not offer much gain in performance-complexity tradeoff. These include when there is lack of suitable conventional iterative algorithms (C1, C2, C4) or when the design problems are highly complicated with constraints and objective function (C3). It can be seen that deep unfolding is not a game changer that will completely replace conventional principle algorithms or black-box ML models. However, the applicability of deep unfolding can be enhanced by overcoming those challenges, as will be discussed next.
XII-B Development of Deep Unfolding Techniques
XII-B1 Coexistence of Deep Unfolding Models, Conventional Solvers, and Black-Box Models

### v3

@@ IX Physical Layer Security
Jamming attacks pose a serious threat to the reliable operation of wireless communication systems [226, 227]. A jamming attacker deliberately transmits interference or noise on the same frequencies as legitimate users. Marti et al. in [228] explored solutions to combat smart jamming in massive multiuser MIMO systems. The authors proposed an optimization-based approach for joint jammer estimation, channel estimation, and data detection, using the BS multiantenna configuration. This method does not require prior knowledge of the jammer’s behavior, making communication more resilient even during sporadic or bursty jamming. The authors have unfolded an iterative algorithm, called as mitigation estimation and detection (MAED), that uses forward-backward splitting optimization to jointly estimate the jammer subspace, user equipment’s channels, and user equipment’s data within a coherence block. The soft-output MAED (SO-MAED), is the deep unfolded network. The trainable set include error precision (to represent noise in symbol estimation), gradient descent step-sizes, scaling factor and momentum weights. SO-MAED improves detection accuracy, especially in high-interference environments, and does not require knowledge of the attack type.
There is limited literature on the application of deep unfolding techniques in physical layer security. One possible reason is that physical layer security strategies are often dynamic and must adapt to adversarial attacks. In contrast, deep unfolding methods typically rely on a structured iterative algorithm which are seldom available in scenarios such as jamming attacks, where the system behavior must adjust rapidly to unpredictable threats.
@@
X Challenges in Developing Deep Unfolding Models and Future Research Scopes
@@
X-A Limitations of Deep Unfolding Techniques
Although deep unfolding has found applications in various design problems, it has limitations. Most of these limitations arise from the fact that deep unfolding is not an independent design methodology; it must be built upon existing conventional optimization frameworks. Consequently, it faces the following challenges in the development and real-world applications:
(C1) Methodological challenges: Deep unfolding methods are inspired by the iterative procedures of conventional optimization algorithms, where each iteration is typically interpreted as a layer in an NN. In traditional wireless transmitter-receiver design and signal processing, optimization problems are often non-convex and highly coupled due to factors such as fractional SINR, constant-modulus constraints (e.g., for analog beamforming), or mixed-integer variables (e.g., in symbol detection or antenna selection). They are typically solved using a two-layer approach. In the first layer, AO or block coordinate ascent/descent is used to decouple variables, breaking the original problem into smaller, more manageable sub-problems. In the second layer, convex optimization techniques such as SCA or the ADMM are employed to solve each sub-problem. While this hierarchical structure enables the use of convex solvers (e.g., CVX), it often introduces auxiliary variables and results in nested loops that must converge at each layer, significantly increasing algorithmic complexity.
Deep unfolding models, on the other hand, aim to transform these iterative procedures into end-to-end trainable architectures. Instead of solving sub-problems sequentially, the unfolding framework typically approximates the entire optimization process in a more unified and parallelizable manner. This shift in design philosophy, from explicitly solving decoupled sub-problems to learning a unified mapping, poses non-trivial challenges, especially when attempting to translate complex multi-stage optimizers with auxiliary variables into a compact, learnable structure. Therefore, while deep unfolding builds upon conventional iterative methods, the methodological differences in structure and implementation represent a major challenge in developing effective unfolding models.
(C2) Unfolding principle algorithms: While deep unfolding typically enhances conventional iterative algorithms, not all can be effectively unfolded. This limitation arises from the core principle of deep unfolding, which structures the iterations of an iterative algorithm as layers within a DNN. Complex operations like matrix decomposition or the use of convex optimization tools such as CVX or YALMIP (with solvers like MOSEK, SeDuMi, etc.) are difficult to implement in a DNN framework. In large-scale systems such as massive MIMO, even matrix multiplications result in high complexity for unfolding models.
@@ X-A Limitations of Deep Unfolding Techniques
(C3) Handling design constraints and objectives: One of the biggest challenges in developing deep unfolding models lies in managing the design constraints formulated in the problem. These constraints require the DNN layers’ outputs to meet specific requirements, such as total power, QoS, or other element-wise constraints. Simple constraints, like power or constant modulus, can often be enforced through normalization [181]. For instance, the $tanh$ activation function can be employed to impose constant-modulus constraints on analog beamforming coefficients or sensing waveforms, similar to black-box DL architectures. However, QoS constraints are significantly more complex. While they resemble activation functions, none of the existing activation functions can adequately manage such intricate requirements. One approach is to incorporate these constraints into the design objective. However, this technique often results in an nonequivalent problem with a more complicated objective function [191].
@@
Not only do complex constraints pose challenges for developing deep unfolding models, but intricate objective functions also present difficulties in training the models. This is because their loss functions are generally formulated based on the objective function. For instance, unfolding architectures like DetNet [49], FSNet [75], and ScNet [229] for symbol detection are trained to minimize the Euclidean distance between the model’s output (symbol vector) and the true transmitted symbol vector. In particular, when labels are unavailable, as in beamforming problems, unfolding networks use the objective function as the loss function for minimization or its negative for maximization problems. Typical deep unfolding techniques [191, 155] require calculating the gradient of the objective function. However, this becomes challenging when the objective function is a complex, high-order, or fractional function of the design variables, especially when those variables are vectors or matrices.
@@
(C4) Sophisticated mappings: While deep unfolding is powerful, certain inference tasks still require black-box DNN models, especially when there is no clear relationship between the input data, the output solution, and performance metrics. For instance, in multi-modal sensing-assisted beam tracking, tasks may involve complex mappings from environmental data, such as sensory information from sensors and cameras, to the optimal beam [230, 231]. In such cases, there is no explicit relationship between environmental images and the optimal beam, and standard communication metrics like throughput or SINR cannot be directly expressed as functions of these images. These problems cannot be formulated as optimization tasks solvable with principle algorithms, making it difficult to develop unfolding models that map environmental information to beam patterns or blockage states.
(C5) Real-world applications: Deep unfolding architectures, while promising in mimicking iterative algorithms through trainable layers, often entail considerable depth and parameter complexity [191]. This imposes severe challenges for real-time inference under strict latency and throughput constraints, especially in systems with multiple and large dimensions like massive MIMO-OFDM. From a hardware perspective, deep unfolding designs are not naturally aligned with the requirements for area-efficient and parallelizable architectures needed in scalable OFDM deployments. For example, the non-uniform and sequential computations common in unfolding detectors can result in inefficient silicon utilization and high power consumption, limiting the ability to instantiate multiple parallel detection cores across subcarriers [232].
Furthermore, in real-world scenarios, the data distribution is highly dynamic because of user mobility, varying propagation environments, and hardware impairments. These variations often violate the assumptions underlying many unfolding-based transmitter and receiver designs, which are typically developed under idealized conditions. However, this assumption breaks down in practical deployments, where channels often exhibit strong line-of-sight components and spatial correlation. As a result, detectors relying on such assumptions can suffer significant performance degradation. To maintain robustness in these conditions, deep unfolding models must be redesigned or trained using data that captures the complexity and variability of realistic, deployment-specific channels.
@@
The above five challenges (C1C5) also outline the scenarios wherein deep unfolding models cannot be built or does not offer much gain in performance-complexity tradeoff. These include when there is lack of suitable conventional iterative algorithms (C1, C2, C4), when the design problems are highly complicated with constraints and objective function (C3), or when the exist imperfection in real-world propagation environments and hardware (C5). It can be seen that deep unfolding is not a game changer that will completely replace conventional principle algorithms or black-box ML models. However, the applicability of deep unfolding can be enhanced by overcoming those challenges, as will be discussed next.
X-B Development of Deep Unfolding Techniques
X-B1 Coexistence of Deep Unfolding Models, Conventional Solvers, and Black-Box Models